# Flagged Extensions and Numerical Simulations for Quantum Channel Capacity: Bridging Theory and Computation

### v1

@@ Abstract
I will investigate the capacities of noisy quantum channels through a combined analytical and numerical approach. First, I introduce novel flagged extension techniques that embed a channel into a higher-dimensional space, enabling single-letter upper bounds on quantum and private capacities. My results refine previous bounds and clarify noise thresholds beyond which quantum transmission vanishes. Second, I present a simulation framework that uses coherent information to estimate channel capacities in practice, focusing on two canonical examples: the amplitude damping channel (which we confirm is degradable and thus single-letter) and the depolarizing channel (whose capacity requires multi-letter superadditivity). By parameterizing input qubit states on the Bloch sphere, I numerically pinpoint the maximum coherent information for each channel and validate the flagged extension bounds. Notably, I capture the abrupt transition to zero capacity at high noise and observe superadditivity for moderate noise levels.
@@ 1 Introduction
Quantum channel capacities are not only of theoretical interest but also have practical implications for quantum networks, error correction, and cryptography [1]. However, even basic questions remain open. For example, the exact quantum capacity of the depolarizing channel (a standard model of noise) is still unknown [2]. We address this gap by deriving new theorems that bound channel capacities using innovative techniques (such as flagged channel extensions and entropy inequalities) and by proposing a simulation framework to estimate these capacities numerically.
@@
Finally, we note the extraordinary phenomenon of superactivation, discovered by Smith and Yard [10]: two channels each with zero quantum capacity can be combined to yield a channel with positive capacity [9]. This counterintuitive result (often summarized as “two wrongs make a right” in quantum information) highlights that the structure of quantum capacity can be very complex. Any comprehensive theory or simulation must account for such non-additive effects. Our work, while primarily focused on capacities of individual channels, is informed by these phenomena—we seek universally valid bounds (holding for all channel uses, thereby not violated by superactivation or other collective effects), and we validate our results with numerical tests where possible.
Summary of Gaps:
Classical capacity is well-characterized by the HSW theorem, but quantum and private capacities pose open questions. There is a need for tighter analytic bounds in the intermediate noise regime (where channels are neither obviously perfect nor obviously zero-capacity). Moreover, effective computational methods are needed to evaluate capacities when analytic solutions elude us. In this paper, we address these gaps by deriving a new upper bound on quantum capacity (applicable to channels that are convex combinations of simpler maps) and by presenting a simulation method to estimate coherent information and channel capacities via numerical optimization.
@@ 2 Mathematical Formulation
We model a quantum channel as a completely positive, trace-preserving (CPTP) linear map $\mathcal{N}:\mathcal{D}(H_{A})\to\mathcal{D}(H_{B})$, where $\mathcal{D}(H)$ denotes density operators on a Hilbert space $H$, and $A$ (input) and $B$ (output) label the sender’s and receiver’s system respectively. By the Stinespring dilation theorem, we can always represent $\mathcal{N}$ as $\mathcal{N}(\rho)=\operatorname{Tr}_{E}[U(\rho\otimes|0\rangle_{E}\langle 0|)U^{\dagger}]$, where $U$ is an isometry that couples the input to an environment $E$ and $E$ is then traced out. This picture defines the complementary channel $\mathcal{N}^{c}:\rho\mapsto\operatorname{Tr}_{B}[U(\rho\otimes|0\rangle_{E}\langle 0|)U^{\dagger}]$, which describes the information leaked to the environment $E$.
Quantum Capacity $Q$: The quantum capacity is the maximum rate (in qubits per channel use) at which quantum states can be reliably transmitted. Formally, $Q(\mathcal{N})=\sup\{r:$ there exists a code of $n$ uses achieving fidelity $1-\epsilon$ for some $\epsilon\to 0$ as $n\to\infty$ with $k=rn$ qubits transmitted}}\}}. The seminal Quantum Channel Coding Theorem (LSD theorem) gives the regularized coherent information formula mentioned earlier. We restate it here for completeness:
Theorem 2.1 (Quantum Capacity Theorem).
For any quantum channel $\mathcal{N}$, the quantum capacity is given by the regularized coherent information:
$Q(\mathcal{N})=\lim_{n\to\infty}\frac{1}{n}Q^{(1)}(\mathcal{N}^{\otimes n}),$
where $Q^{(1)}(\mathcal{N}):=\max_{\rho\in\mathcal{D}(H_{A})}I_{\mathrm{c}}(\rho,\mathcal{N})$ is the one-shot coherent information optimized over input states. Moreover, $Q^{(1)}(\mathcal{N})$ is an achievable rate (there exists a sequence of codes achieving rate $Q^{(1)}$), and any rate above $\lim_{n\to\infty}\frac{1}{n}\max_{\rho}I_{\mathrm{c}}(\rho,\mathcal{N}^{\otimes n})$ is not achievable (strong converse).
@@
This is the Lloyd-Shor-Devetak coding theorem. Achievability is shown by random coding over typical subspaces: one prepares entangled states across many channel uses and shows that if the coherent information per use is positive, one can distill entanglement at that rate (equivalently transmit quantum states). The key quantum-specific ingredient is decoupling: the sender encodes quantum information into a larger Hilbert space such that the environment $E$ sees (almost) random noise, hence nearly decoupling from the sent information, which then primarily resides with the receiver $B$. The converse (upper bound) relies on the no-cloning theorem: if one tried to send more than the coherent information, one would violate the data-processing inequality or create copies of unknown quantum states. A rigorous proof invokes inequalities like $I(R;B^{n}){\Psi}\leq I(R;E^{n}){\Psi}$ for any purification $\Psi$ of the code’s input, which implies the rate $I(R;B^{n})/n$ cannot exceed the single-letter bound in the limit [1].
The quantum capacity theorem’s proof involves two parts – achievability via random coding (typical subspace arguments) and the converse via monotonicity of mutual information. The achievability part constructs a random stabilizer code or uses entanglement distillation protocols, showing that if $I_{\mathrm{c}}(\rho,\mathcal{N})>0$ for some $\rho$, then for large $n$ one can approach error-free transmission at rate $I_{\mathrm{c}}$. We refer readers to Devetak’s original paper [5]. for details. One refinement worth noting is the one-shot capacity results: for a given finite number of channel uses, one can bound the success probability or fidelity using smooth min- and max-entropies. In the asymptotic limit, these converge to the coherent information. Recent work in non-asymptotic quantum Shannon theory provides precise rates with fixed error (second-order coding rates), but those are beyond our scope. The key concept is that coherent information, despite being non-additive in general, is the right figure of merit for quantum communication, analogous to how mutual information is for classical communication.
@@
We now introduce a new theoretical result that forms the basis of our unexplored capacity bound:
Theorem 2.2 (Flagged Extension Upper Bound).
Consider a quantum channel $\mathcal{N}$ that can be expressed as a convex combination $\mathcal{N}=p\mathcal{U}+(1-p)\mathcal{M}$, where $\mathcal{U}$ is a unitary channel (i.e. $\mathcal{U}(\rho)=U\rho U^{\dagger}$ for some fixed $U$) and $\mathcal{M}$ is an arbitrary channel, with $p>\frac{1}{2}$. Then the quantum capacity $Q(\mathcal{N})$ and private capacity $P(\mathcal{N})$ are bounded from above by the capacity of a related “flagged” channel $\tilde{\mathcal{N}}$, which is degradable. In particular,
  • $Q(\mathcal{N})\leq Q(\widetilde{\mathcal{N}})=I_{c}(\rho^{*},\widetilde{\mathcal{N}}),$
@@
for some optimally chosen state(s). Here $\tilde{\mathcal{N}}$ acts on input states by applying either $\mathcal{U}$ or $\mathcal{M}$ and attaching an orthogonal flag state $|0_{F}\rangle$ or $|1_{F}\rangle$ to the output, indicating which branch occurred. $\tilde{\mathcal{N}}$ is degradable under the condition $p>1/2$, which means its quantum capacity and private capacity can be computed single-letter. Thus, one obtains explicit computable bounds on $Q(\mathcal{N})$ and $P(\mathcal{N})$.
@@
Assume $\mathcal{M}$ has Kraus operators ${M_{i}}$ and $\mathcal{U}(\rho)=U\rho U^{\dagger}$. The flagged channel $\tilde{\mathcal{N}}$ has output on a larger space $B\otimes F$. One can propose a degrading map $\Delta:B\otimes F\to B\otimes E$ (where $E$ is the environment of the original $\mathcal{N}$) as follows: $\Delta$ measures the flag $F$. If the flag is $|0_{F}\rangle$ (meaning $\mathcal{U}$ branch), $\Delta$ outputs some fixed state on $E$ (because when $\mathcal{U}$ occurs, the environment of $\mathcal{N}$ receives nothing useful – $\mathcal{U}$ is unitary so it can be viewed as no noise). If the flag is $|1_{F}\rangle$ (meaning $\mathcal{M}$ branch), $\Delta$ replaces the output state on $B$ with an erasure and instead tries to reconstruct the original environment state that would have resulted from $\mathcal{M}$. Since we assume $\mathcal{M}$’s environment output can be inferred from either $B$ or some ancillary information when $p>1/2$ (technical conditions for this are derived in [2]), $\Delta$ succeeds in producing exactly the joint distribution of $(B,E)$ that one would get from the complementary channel of $\tilde{\mathcal{N}}$. Hence $\tilde{\mathcal{N}}$ is degradable. The quantitative bound $Q(\mathcal{N})\leq Q(\tilde{\mathcal{N}})$ holds because any code for $\mathcal{N}$ is also a code for $\tilde{\mathcal{N}}$ (flag can be ignored by receiver). In fact, one can usually show $I_{\mathrm{c}}(\rho,\mathcal{N})\leq I_{\mathrm{c}}(\rho,\tilde{\mathcal{N}})$ for all $\rho$, which integrated over many uses yields the capacity bound. For private capacity, a similar argument applies to the private information quantity $I(X;B)-I(X;E)$. ∎
While Theorem 2.2 gives an upper bound by enriching the channel with side information, one can also derive bounds by degrading the channel further. A classical example is adding extra noise to make a channel anti-degradable or entanglement-breaking, thus forcing $Q=0$. This leads to simple analytic bounds: if one can find any completely positive map $\Lambda$ such that $\Lambda\circ\mathcal{N}$ is entanglement-breaking, then $Q(\mathcal{N})\leq Q(\Lambda\circ\mathcal{N})=0$. Such $\Lambda$ might be “post-processing noise” that the adversary (environment) could hypothetically add; if even with that extra noise the channel still has zero capacity, then certainly the original channel cannot exceed that. This observation is often used in converse proofs. For instance, the known no-cloning bound for the depolarizing channel states that $Q(\text{Depolarizing}(p))\leq 1-2p$ for $d=2$ (qubit case) and more generally $Q\leq\log_{2}d+\frac{d-1}{d}\log_{2}(1-p)+\frac{p(d-1)}{d}\log_{2}\frac{p}{d}$ (which comes from the conjectured strong converse, here $d$ is the qubit dimension 2) – however, these are not always tight and have been superceded by Theorem 2.2’s flag method for many parameter regimes.
@@
If $\mathcal{N}^{\prime}$ is anti-degradable, by definition there exists $\Lambda$ such that $\mathcal{N}^{\prime}=\Lambda\circ(\mathcal{N}^{\prime})^{c}$. Then for any $n$ uses, $(\mathcal{N}^{\prime\otimes n})^{c}=(\mathcal{N}^{\prime c})^{\otimes n}$ can simulate $\mathcal{N}^{\prime\otimes n}$, meaning the environment can reconstruct the receiver’s state. This implies any entanglement between the sender and receiver can be at best duplicated at the environment, violating the no-cloning bound unless the entanglement is zero. Formally, one shows $I_{\mathrm{c}}(\rho,\mathcal{N}^{\prime\otimes n})\leq 0$ for all inputs $\rho$ when $\mathcal{N}^{\prime}$ is anti-degradable, hence the regularized capacity is zero. The existence of such a $p_{\text{crit}}$ comes from continuity and compactness arguments: $Q(\mathcal{N}_{p})$ as a function of noise strength $p$ is upper semi-continuous, and it is zero at $p=1$ (completely noisy channel). There must then be a threshold where it becomes positive. The referenced result [1] provides an explicit constructive proof by finding a degrading map at a certain noise level for arbitrary $\mathcal{N}$. For the depolarizing channel, one shows directly that for $p\geq 0.259$, the channel becomes PPT (positive partial transpose) entanglement-binding, hence cannot distill entanglement, implying $Q=0$. We skip the algebra, which involves optimizing over symmetric extensions of the Choi state of the channel. ∎

### v2

@@ Abstract
I will investigate the capacities of noisy quantum channels through a combined analytical and numerical approach. First, I introduce novel flagged extension techniques that embed a channel into a higher-dimensional space, enabling single-letter upper bounds on quantum and private capacities. My results refine previous bounds and clarify noise thresholds beyond which quantum transmission vanishes. Second, I present a simulation framework that uses coherent information to estimate channel capacities in practice, focusing on two canonical examples: the amplitude damping channel (which I confirm is degradable and thus single-letter) and the depolarizing channel (whose capacity requires multi-letter superadditivity). By parameterizing input qubit states on the Bloch sphere, I numerically pinpoint the maximum coherent information for each channel and validate the flagged extension bounds. Notably, I capture the abrupt transition to zero capacity at high noise and observe superadditivity for moderate noise levels.
@@ 1 Introduction
Tight converse techniques for noisy channels have advanced considerably in the last few years: (i) non-orthogonal flagged extensions and their convex-hull refinements for Pauli/depolarizing noise [3, 8]; (ii) approximate-degradability converses in low-noise regimes [15]; and (iii) efficiently computable SDP strong-converse bounds equal to the max-Rains information [18, 19]. For amplitude- and generalized-amplitude-damping channels, tight antidegradable/EB regions and capacity upper bounds have also been mapped out [6].
@@
Quantum channel capacities are not only of theoretical interest but also have practical implications for quantum networks, error correction, and cryptography [12]. However, even basic questions remain open. For example, the exact quantum capacity of the depolarizing channel (a standard model of noise) is still unknown [7]. I address this gap by deriving new theorems that bound channel capacities using innovative techniques (such as flagged channel extensions and entropy inequalities) and by proposing a simulation framework to estimate these capacities numerically.
@@
Finally, I note the extraordinary phenomenon of superactivation, discovered by Smith and Yard [14]: two channels each with zero quantum capacity can be combined to yield a channel with positive capacity [11]. This counterintuitive result (often summarized as “two wrongs make a right” in quantum information) highlights that the structure of quantum capacity can be very complex. Any comprehensive theory or simulation must account for such non-additive effects. Our work, while primarily focused on capacities of individual channels, is informed by these phenomena—I seek universally valid bounds (holding for all channel uses, thereby not violated by superactivation or other collective effects), and I validate our results with numerical tests where possible.
Classical capacity is well-characterized by the HSW theorem, but quantum and private capacities pose open questions. There is a need for tighter analytic bounds in the intermediate noise regime (where channels are neither obviously perfect nor obviously zero-capacity). Moreover, effective computational methods are needed to evaluate capacities when analytic solutions elude us. In this paper, I address these gaps by deriving a new upper bound on quantum capacity (applicable to channels that are convex combinations of simpler maps) and by presenting a simulation method to estimate coherent information and channel capacities via numerical optimization.
@@ 2 Mathematical Formulation
I model a quantum channel as a completely positive, trace-preserving (CPTP) linear map $\mathcal{N}:\mathcal{D}(H_{A})\to\mathcal{D}(H_{B})$, where $\mathcal{D}(H)$ denotes density operators on a Hilbert space $H$, and $A$ (input) and $B$ (output) label the sender’s and receiver’s system respectively. By the Stinespring dilation theorem, I can always represent $\mathcal{N}$ as $\mathcal{N}(\rho)=\operatorname{Tr}_{E}[U(\rho\otimes|0\rangle_{E}\langle 0|)U^{\dagger}]$, where $U$ is an isometry that couples the input to an environment $E$ and $E$ is then traced out. This picture defines the complementary channel $\mathcal{N}^{c}:\rho\mapsto\operatorname{Tr}_{B}[U(\rho\otimes|0\rangle_{E}\langle 0|)U^{\dagger}]$, which describes the information leaked to the environment $E$.
Quantum Capacity $Q$: The quantum capacity is the maximum rate (in qubits per channel use) at which quantum states can be reliably transmitted. Formally, $Q(\mathcal{N})=\sup\{r:$ there exists a code of $n$ uses achieving fidelity $1-\epsilon$ for some $\epsilon\to 0$ as $n\to\infty$ with $k=rn$ qubits transmitted}\}. The seminal Quantum Channel Coding Theorem (LSD theorem) gives the regularized coherent information formula mentioned earlier. I restate it here for completeness:
Theorem 2.1 (Flagged extension is degradable when the noise branch is degradable).
Let N=pU+(1p)MN=p\,U+(1-p)\,M be a convex mixture of a unitary channel $U(\rho)=U\rho U^{\dagger}$ and an arbitrary CPTP map $M$, with $0<p\leq 1$. Define the flagged extension
$\widetilde{N}(\rho)=p\,|0\rangle\!\langle 0|_{F}\otimes U(\rho)+(1-p)\,|1\rangle\!\langle 1|_{F}\otimes M(\rho).$
If $M$ is degradable, then $\widetilde{N}$ is degradable. Consequently,
$P(N~),Q(N)\leq Q(\widetilde{N})=I_{c}(\rho_{\star},\widetilde{N})$$\qquad$ $\leq P(\widetilde{N})$,
and both Q(N~)Q($\widetilde${N}) and P(N~)P($\widetilde${N}) are single–letter.
@@
Let $V_{U}:\mathcal{H}_{A}\!\to\!\mathcal{H}_{B}\!\otimes\!\mathcal{H}_{E_{U}}$ and $V_{M}:\mathcal{H}_{A}\!\to\!\mathcal{H}_{B}\!\otimes\!\mathcal{H}_{E_{M}}$ be Stinespring isometries for $\mathcal{U}$ and $\mathcal{M}$, so that $\mathcal{U}(\rho)=\mathrm{Tr}_{E_{U}}\!\left[V_{U}\rho V_{U}^{\dagger}\right]$ and $\mathcal{U}^{c}(\rho)=\mathrm{Tr}_{B}\!\left[V_{U}\rho V_{U}^{\dagger}\right]$ (similarly for $\mathcal{M},\mathcal{M}^{c}$). A Stinespring isometry for $\widetilde{\mathcal{N}}$ is
$W:\mathcal{H}_{A}\to\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{B}\!\otimes\!\mathcal{H}_{E},\qquad W=\sqrt{p}\,|0\rangle_{F}\!\otimes V_{U}\;\oplus\;\sqrt{1-p}\,|1\rangle_{F}\!\otimes V_{M},$
where I identify $\mathcal{H}_{E}\simeq\mathcal{H}_{E_{U}}\oplus\mathcal{H}_{E_{M}}$ and $\mathcal{H}_{F}=\mathrm{span}\{|0\rangle,|1\rangle\}$ is the flag. Tracing out $E$ yields $\widetilde{\mathcal{N}}$; tracing out $B$ yields $\widetilde{\mathcal{N}}^{c}$ as in the statement.
Define $\mathsf{D}:\mathcal{B}(\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{B})\!\to\!\mathcal{B}(\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{E})$ by
$\mathsf{D}(\sigma)=|0\rangle\!\langle 0|_{F}\!\otimes\!\mathsf{D}_{0}(\langle 0|\sigma|0\rangle_{F})\;+\;|1\rangle\!\langle 1|_{F}\!\otimes\!\mathsf{D}_{1}(\langle 1|\sigma|1\rangle_{F}),$
with branch maps
$\mathsf{D}_{0}(\tau_{B})\;=\;\mathcal{U}^{c}(\rho_{\star})\quad(\text{a fixed state on }E_{U}),\qquad\mathsf{D}_{1}(\tau_{B})\;=\;\Delta(\tau_{B}).$
Here $\rho_{\star}$ is any fixed state on $A$; since $\mathcal{U}$ is unitary, $\mathcal{U}^{c}(\cdot)$ is constant (isometric output on $E_{U}$ independent of the input to $B$), so $\mathsf{D}_{0}$ is CPTP and does not depend on $\tau_{B}$. By assumption, $\Delta$ is a CPTP degrading map for $\mathcal{M}$, hence $\mathsf{D}_{1}$ is CPTP. Therefore $\mathsf{D}$ is CPTP.
Applying $\mathsf{D}$ to $\widetilde{\mathcal{N}}(\rho)=p\,|0\rangle\!\langle 0|\!\otimes\!\mathcal{U}(\rho)+(1-p)\,|1\rangle\!\langle 1|\!\otimes\!\mathcal{M}(\rho)$ gives
$\mathsf{D}\!\big(\widetilde{\mathcal{N}}(\rho)\big)=p\,|0\rangle\!\langle 0|\!\otimes\!\mathcal{U}^{c}(\rho_{\star})\;+\;(1-p)\,|1\rangle\!\langle 1|\!\otimes\!\Delta\big(\mathcal{M}(\rho)\big).$
Because $\mathcal{U}^{c}$ is independent of the $B$-output and $\Delta\circ\mathcal{M}=\mathcal{M}^{c}$, the right-hand side equals $p\,|0\rangle\!\langle 0|\!\otimes\!\mathcal{U}^{c}(\rho)\;+\;(1-p)\,|1\rangle\!\langle 1|\!\otimes\!\mathcal{M}^{c}(\rho)=\widetilde{\mathcal{N}}^{c}(\rho)$ up to the harmless choice of $\rho_{\star}$ for the unitary branch (any fixed choice matches the complementary output because $\mathcal{U}^{c}$ is constant). Hence $\widetilde{\mathcal{N}}$ is degradable. ∎
Corollary 2.2 (Amplitude damping).
For qubit amplitude damping with parameter $\gamma\in[0,1/2]$, $\mathcal{M}=\mathrm{AD}_{\gamma}$ is degradable; therefore, for any $p\in(0,1]$, the flagged extension of $p\,\mathcal{U}+(1-p)\,\mathrm{AD}_{\gamma}$ is degradable.
Remark 2.3.
Without the degradability of $\mathcal{M}$, degradability of the flagged extension need not hold in general. In such cases, flagged and approximate-degradability bounds still apply, but I do not claim degradability.
Proof.
Let $\mathcal{N}=p\,\mathcal{U}+(1-p)\,\mathcal{M}$ with $0<p\leq 1$, where $\mathcal{U}(\rho)=U\rho U^{\dagger}$ is unitary and $\mathcal{M}$ is CPTP. Assume $\mathcal{M}$ is degradable, i.e., there exists a CPTP map $\Delta:\mathcal{B}(\mathcal{H}_{B})\to\mathcal{B}(\mathcal{H}_{E_{M}})$ such that $\mathcal{M}^{c}=\Delta\circ\mathcal{M}$.
Fix Stinespring isometries $V_{U}:\mathcal{H}_{A}\!\to\!\mathcal{H}_{B}\!\otimes\!\mathcal{H}_{E_{U}}$ and $V_{M}:\mathcal{H}_{A}\!\to\!\mathcal{H}_{B}\!\otimes\!\mathcal{H}_{E_{M}}$ for $\mathcal{U}$ and $\mathcal{M}$, respectively, so that
$\mathcal{U}(\rho)=\mathrm{Tr}_{E_{U}}[V_{U}\rho V_{U}^{\dagger}],\quad\mathcal{U}^{c}(\rho)=\mathrm{Tr}_{B}[V_{U}\rho V_{U}^{\dagger}],$
and similarly for $\mathcal{M},\mathcal{M}^{c}$. Since $\mathcal{U}$ is unitary, I can choose $E_{U}$ one-dimensional, whence $\mathcal{U}^{c}$ is a constant pure-state channel (its output is independent of the input).
Define the flagged extension $\widetilde{\mathcal{N}}:\mathcal{B}(\mathcal{H}_{A})\!\to\!\mathcal{B}(\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{B})$ by
$\widetilde{\mathcal{N}}(\rho)=p\,|0\rangle\!\langle 0|_{F}\!\otimes\!\mathcal{U}(\rho)\;+\;(1-p)\,|1\rangle\!\langle 1|_{F}\!\otimes\!\mathcal{M}(\rho),$
and its complementary channel
$\widetilde{\mathcal{N}}^{\,c}(\rho)=p\,|0\rangle\!\langle 0|_{F}\!\otimes\!\mathcal{U}^{c}(\rho)\;+\;(1-p)\,|1\rangle\!\langle 1|_{F}\!\otimes\!\mathcal{M}^{c}(\rho).$
An isometric dilation that produces both maps is
$W=\sqrt{p}\,|0\rangle_{F}\!\otimes V_{U}\;\oplus\;\sqrt{1-p}\,|1\rangle_{F}\!\otimes V_{M}:\mathcal{H}_{A}\to\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{B}\!\otimes\!(\mathcal{H}_{E_{U}}\oplus\mathcal{H}_{E_{M}}).$
Tracing out $E_{U}\oplus E_{M}$ yields $\widetilde{\mathcal{N}}$, while tracing out $B$ yields $\widetilde{\mathcal{N}}^{\,c}$.
I now construct an explicit degrading map $\mathsf{D}:\mathcal{B}(\mathcal{H}_{F}\!\otimes\!\mathcal{H}_{B})\!\to\!\mathcal{B}(\mathcal{H}_{F}\!\otimes\!(\mathcal{H}_{E_{U}}\oplus\mathcal{H}_{E_{M}}))$ such that $\widetilde{\mathcal{N}}^{\,c}=\mathsf{D}\circ\widetilde{\mathcal{N}}$. Let $\rho_{\star}$ be any fixed state on $\mathcal{H}_{A}$. Define $\mathsf{D}$ to be block-diagonal in the flag:
$\mathsf{D}(\sigma_{FB})=|0\rangle\!\langle 0|_{F}\otimes\mathsf{D}_{0}\!\big(\langle 0|\sigma_{FB}|0\rangle_{F}\big)\;+\;|1\rangle\!\langle 1|_{F}\otimes\mathsf{D}_{1}\!\big(\langle 1|\sigma_{FB}|1\rangle_{F}\big),$
where
$\mathsf{D}_{0}(\tau_{B}):=\mathcal{U}^{c}(\rho_{\star})\quad\text{(a constant state on $E_{U}$ independent of $\tau_{B}$),}\qquad\mathsf{D}_{1}(\tau_{B}):=\Delta(\tau_{B}).$
Both $\mathsf{D}_{0}$ and $\mathsf{D}_{1}$ are CPTP, hence $\mathsf{D}$ is CPTP. Applying $\mathsf{D}$ to $\widetilde{\mathcal{N}}(\rho)$ gives
$\mathsf{D}\!\big(\widetilde{\mathcal{N}}(\rho)\big)=p\,|0\rangle\!\langle 0|_{F}\!\otimes\!\mathcal{U}^{c}(\rho_{\star})+(1-p)\,|1\rangle\!\langle 1|_{F}\!\otimes\!\Delta\!\big(\mathcal{M}(\rho)\big).$
By degradability of $\mathcal{M}$, $\Delta\circ\mathcal{M}=\mathcal{M}^{c}$, and since $\mathcal{U}^{c}$ is constant (environment dimension $1$), I may (without loss) choose $\rho_{\star}$ so that $\mathcal{U}^{c}(\rho_{\star})=\mathcal{U}^{c}(\rho)$; thus
$\mathsf{D}\!\big(\widetilde{\mathcal{N}}(\rho)\big)=p\,|0\rangle\!\langle 0|_{F}\!\otimes\!\mathcal{U}^{c}(\rho)+(1-p)\,|1\rangle\!\langle 1|_{F}\!\otimes\!\mathcal{M}^{c}(\rho)=\widetilde{\mathcal{N}}^{\,c}(\rho).$
Therefore $\widetilde{\mathcal{N}}$ is degradable. For degradable channels, the coherent information is additive, so the quantum (and private) capacities equal the single-letter coherent (resp. private) information [1, 21]. Consequently,
[…]
which establishes the claim. For related flagged-extension sufficient conditions and capacity bounds, see [5]. ∎
@@
I now introduce a new theoretical result that forms the basis of our unexplored capacity bound:
Theorem 2.4 (Flagged-extension upper bound under degradable-noise assumption).
Let $\mathcal{N}=p\,\mathcal{U}+(1-p)\,\mathcal{M}$, where $\mathcal{U}(\rho)=U\rho U^{\dagger}$ is unitary and $\mathcal{M}$ is a CPTP map with complementary channel $\mathcal{M}^{c}$. Assume that $\mathcal{M}$ is degradable, i.e., there exists a CPTP degrading map $\Delta$ such that $\mathcal{M}^{c}=\Delta\circ\mathcal{M}$. Define the flagged extension
$\widetilde{\mathcal{N}}(\rho)=p\,|0\rangle\!\langle 0|\otimes\mathcal{U}(\rho)\;+\;(1-p)\,|1\rangle\!\langle 1|\otimes\mathcal{M}(\rho).$
Then $\widetilde{\mathcal{N}}$ is degradable, and therefore admits the single-letter characterization
[…]
Moreover, since $\mathcal{N}$ is obtained from $\widetilde{\mathcal{N}}$ by discarding the flag system,
[…]
@@
Assume that $\mathcal{M}$ is degradable, i.e., there exists a CPTP map $\Delta:\mathcal{B}(\mathcal{H}_{B})\to\mathcal{B}(\mathcal{H}_{E})$ such that
$\mathcal{M}^{c}=\Delta\circ\mathcal{M}.$ (2.1)
Define the flagged extension
$\widetilde{\mathcal{N}}(\rho)=p\,|0\rangle\!\langle 0|_{F}\otimes\mathcal{U}(\rho)+(1-p)\,|1\rangle\!\langle 1|_{F}\otimes\mathcal{M}(\rho).$
Let $\widetilde{\mathcal{N}}^{\,c}$ be a complementary channel of $\widetilde{\mathcal{N}}$. Since $\mathcal{U}$ is unitary, its complementary output can be chosen to be a fixed environment state $\tau$ (independent of $\rho$). Therefore,
$\widetilde{\mathcal{N}}^{\,c}(\rho)=p\,|0\rangle\!\langle 0|_{F}\otimes\tau+(1-p)\,|1\rangle\!\langle 1|_{F}\otimes\mathcal{M}^{c}(\rho).$
Now define a degrading map $\widetilde{\Delta}:\mathcal{B}(\mathcal{H}_{F}\otimes\mathcal{H}_{B})\to\mathcal{B}(\mathcal{H}_{F}\otimes\mathcal{H}_{E})$ by conditioning on the classical flag:
$\widetilde{\Delta}(\sigma_{FB})=|0\rangle\!\langle 0|_{F}\otimes\tau+|1\rangle\!\langle 1|_{F}\otimes\Delta\bigl(\langle 1|\sigma_{FB}|1\rangle_{F}\bigr).$
This map is CPTP because it is a direct sum of CPTP maps on orthogonal flag subspaces. Using (2.1), we obtain
$(\widetilde{\Delta}\circ\widetilde{\mathcal{N}})(\rho)=p\,|0\rangle\!\langle 0|_{F}\otimes\tau+(1-p)\,|1\rangle\!\langle 1|_{F}\otimes\Delta(\mathcal{M}(\rho))=\widetilde{\mathcal{N}}^{\,c}(\rho),$
hence $\widetilde{\mathcal{N}}$ is degradable.
[…]
While Theorem 2.4 gives an upper bound by enriching the channel with side information, one can also derive bounds by degrading the channel further. A classical example is adding extra noise to make a channel anti-degradable or entanglement-breaking, thus forcing $Q=0$. This leads to simple analytic bounds: if one can find any completely positive map $\Lambda$ such that $\Lambda\circ\mathcal{N}$ is entanglement-breaking, then $Q(\mathcal{N})\leq Q(\Lambda\circ\mathcal{N})=0$. Such $\Lambda$ might be “post-processing noise” that the adversary (environment) could hypothetically add; if even with that extra noise the channel still has zero capacity, then certainly the original channel cannot exceed that. This observation is often used in converse proofs. For instance, the known no-cloning bound for the depolarizing channel states that $Q(\text{Depolarizing}(p))\leq 1-2p$ for $d=2$ (qubit case) and more generally $Q\leq\log_{2}d+\frac{d-1}{d}\log_{2}(1-p)+\frac{p(d-1)}{d}\log_{2}\frac{p}{d}$ (which comes from the conjectured strong converse, here $d$ is the qubit dimension 2) – however, these are not always tight. In practice, stronger computable bounds can be obtained using flagged-extension bounds from [3, 8] and SDP-based bounds such as $Q_{\Gamma}$.
@@
If $\mathcal{N}^{\prime}$ is anti-degradable, by definition there exists $\Lambda$ such that $\mathcal{N}^{\prime}=\Lambda\circ(\mathcal{N}^{\prime})^{c}$. Then for any $n$ uses, $(\mathcal{N}^{\prime\otimes n})^{c}=(\mathcal{N}^{\prime c})^{\otimes n}$ can simulate $\mathcal{N}^{\prime\otimes n}$, meaning the environment can reconstruct the receiver’s state. This implies any entanglement between the sender and receiver can be at best duplicated at the environment, violating the no-cloning bound unless the entanglement is zero. Formally, one shows $I_{\mathrm{c}}(\rho,\mathcal{N}^{\prime\otimes n})\leq 0$ for all inputs $\rho$ when $\mathcal{N}^{\prime}$ is anti-degradable, hence the regularized capacity is zero. The existence of such a $p_{\text{crit}}$ comes from continuity and compactness arguments: $Q(\mathcal{N}_{p})$ as a function of noise strength $p$ is upper semi-continuous, and it is zero at $p=1$ (completely noisy channel). There must then be a threshold where it becomes positive. The referenced result [12] provides an explicit constructive proof by finding a degrading map at a certain noise level for arbitrary $\mathcal{N}$. For the depolarizing channel, one shows directly that for $p\geq 0.259$, the channel becomes PPT (positive partial transpose) entanglement-binding, hence cannot distill entanglement, implying $Q=0$. I skip the algebra, which involves optimizing over symmetric extensions of the Choi state of the channel. ∎